# Constant-Time Motion Planning with Manipulation Behaviors

### v1

@@ Abstract
Recent progress in contact-rich robotic manipulation has been striking, yet most deployed systems remain confined to simple, scripted routines. One of the key barriers is the lack of motion planning algorithms that can provide verifiable guarantees for safety, efficiency and reliability. To address this, a family of algorithms called Constant-Time Motion Planning (CTMP) was introduced, which leverages a preprocessing phase to enable collision-free motion queries in a fixed, user-specified time budget (e.g., 10 milliseconds). However, existing CTMP methods do not explicitly incorporate the manipulation behaviors essential for object handling. To bridge this gap, we introduce the Behavioral Constant-Time Motion Planner (B-CTMP), an algorithm that extends CTMP to solve a broad class of two-step manipulation tasks: (1) a collision-free motion to a behavior initiation state, followed by (2) execution of a manipulation behavior (such as grasping or insertion) to reach the goal. By precomputing compact data structures, B-CTMP guarantees constant-time query in mere milliseconds while ensuring completeness and successful task execution over a specified set of states. We evaluate B-CTMP on two canonical manipulation tasks in simulation, shelf picking and plug insertion,and demonstrate its effectiveness on a real robot. Our results show that B-CTMP unifies collision-free planning and object manipulation within a single constant-time framework, providing provable guarantees of speed and success for manipulation in semi-structured environments.
@@ I Introduction
Despite remarkable advances in robotic manipulation research, a persistent gap remains between the capabilities demonstrated in research laboratories and the requirements for real-world industrial deployment. This gap is especially pronounced in dynamic, semi-structured environments such as warehouse shelf picking (e.g., Amazon fulfillment centers [1]), bin picking in logistics, and precision assembly in manufacturing. A central challenge is the lack of verifiable guarantees on system performance—particularly safety, efficiency, and predictability—which are essential for practitioners in safety-critical and high-throughput applications.
@@ II-A Motion Planning with Preprocessing
To address these limitations, recent approaches focus on constructing collision-free regions in the configuration space rather than sampling discrete states, making motion within these regions computationally efficient during online planning. A prominent example of this region-based approach is Planning in Graphs of Convex Sets [8, 9, 10, 11, 12] which decomposes the configuration space offline into collision-free convex sets [13, 14], enabling fast generation of smooth motion plans. These algorithms offer several advantages: they enable smooth trajectory generation through convex optimization and can handle complex geometric constraints naturally. However, they do not guarantee time bounds on online planning duration.
@@ III Preliminary
To build towards our aim, we first define the concept of coverage to capture behavior feasibility from specific robot states.
@@ IV-A Preprocessing Phase
Instead, we exploit a key insight: manipulation behaviors often exhibit spatial locality, where a single initiation state can cover multiple object configurations within a spatial region. For example, when grasping objects from a shelf, one pre-grasp pose (initiation state) may enable successful grasping of any object within a certain shelf region, as the behavior can adapt to the specific object pose during execution. Hence, leveraging this coverage property of initiation states, we can strategically select a reduced set of feasible initiation states $\widetilde{\mathcal{S}}\subseteq\displaystyle{\bigcup_{w_{g}\in\mathcal{G}}I_{\sigma}(w_{g})}$ such that:
$\bigcup_{s_{i}\in\widetilde{\mathcal{S}}}n_{i}(s_{i})\supseteq\mathcal{G},$
which provides complete coverage over all desired goal states while requiring significantly fewer paths from $s_{home}$.
@@
Figure 2: Overview of the preprocessing structure. The planner computes a reduced set of feasible initiation states $\widetilde{\mathcal{S}}$, each connected to the robot’s home state $s_{home}$ by a precomputed path. Each $s_{i}\in\widetilde{\mathcal{S}}$ defines a neighborhood $n_{i}(s_{i})$ within the region of interest $\mathcal{G}$ of object states $w$ that can be reached through behavior execution. This compression is captured by attractor tuples, each representing an initiation state, its reachable neighborhood, and the stored path that enables efficient plan retrieval during online execution.
@@
Once the algorithm encounters an object state that cannot be reached by the behavior rollout from $s_{i}$, that state is recorded and subsequently used to seed coverage for a different local region of interest. Once all the candidate neighborhoods are expanded, the largest neighborhood (measured in terms of the maximum distance $r$ between the attractor object state $w_{attr}$ and the closest frontier state) is selected to get added to the stored library, and the initiation state corresponding to the neighborhood is chosen as the attractor initiation state $s_{i,attr}$.
@@ V Experiments
We evaluate our B-CTMP on two canonical manipulation tasks across distinct simulation environments–shelf picking and plug insertion, and demonstrate its effectiveness on a physical UR10e robot in the accompanying video.
@@ V-A1 Shelf Grasping
For our first task, we consider a shelf-picking scenario commonly encountered in industrial warehouse automation, where a robotic manipulator must retrieve objects from structured storage environments. The manipulation task involves grasping a target object whose pose is specified in the world coordinate frame by $w_{g}\in\mathrm{SE}(3)$.
Grasp Behavior: The grasp behavior $\sigma_{grasp}$ depends on two key points, the pre-grasp pose (an initiate state) from which we execute the policy to reach the second key point, and the grasp pose, where we activate a closure sequence to grasp the object. The behavior policy $\pi_{\sigma_{grasp}}$ uses Jacobian control to gradually and smoothly move the end-effector pose through differential kinematics between these points, enabling grasp execution.
Behavior Implementation: The GetInitStates function for the grasp behavior outputs a subset of valid pre-grasp poses (initiation states), which are computed as follows. First, we compute candidate grasp poses with valid IK solutions using the target object’s point cloud. Then, Local surface normals are used to calculate antipodal scores that quantify grasp stability by measuring alignment between gripper fingers and object surface. The top $K$ grasps ranked by antipodal score are selected, and the corresponding pre-grasp poses are computed by applying a fixed retraction transformation along each grasp’s approach direction.
B-CTMP Integration: B-CTMP leverages the behavior during offline preprocessing to compute a compact set of representative initiation states with their corresponding neighborhoods. During this process, the system discovers candidate states and selects those with the largest coverage areas to serve as tuple-attractors in the final library, creating a small set of representative pre-grasp poses capable of handling many different target object configurations.
@@ V-A2 Charger Insertion
The second task we consider is precision charger insertion with perception, commonly required in automated charging systems and electronic assembly. This task requires inserting a charger connector into a target port located at $w_{g}\in SE(3)$, and represents dual challenges due to strict geometric constraints for successful insertion and uncertainty from perception.
Insertion Behavior: In this case, since the geometry of the port uniquely dictates a single insertion pose, all the behavior needs is the initiation state from which to rollout the policy. However, due to the high-precision required in the task and the perception uncertainty, we rely on visual-based Jacobian control for our policy $\pi_{\sigma_{insert}}$.
Behavior Implementation: We directly compute the insertion pose for each port location and apply a predefined transformation to obtain the corresponding pre-insert initiation state, which is output by the GetInitStates function. To ensure robust execution under perception uncertainty, we impose a manipulability constraint during policy execution. We define the minimum manipulability radius as $r_{\min}(q)=\min_{i}\sqrt{\lambda_{i}\!\big(J(q)J(q)^{\mathsf{T}}\big)}$ where the manipulability ellipsoid radii are derived from the Jacobian eigenvalues.
B-CTMP Integration: B-CTMP’s offline preprocessing step creates a compact set of initiation states that can cover multiple port locations while maintaining the ability to handle the perception uncertainty.
It does so by imposing the manipulability constraint during behavior rollouts during preprocessing, marking the rollout as infeasible if $r_{\min}(q)<\epsilon$, where $\epsilon$ represents the perception noise. This ensures that selected attractor pre-insert poses can maintain adequate dexterity for corrective motions across their neighborhoods.
B-CTMP performs all pose discovery processes (grasp pose computation and pre-insertion pose calculation) offline to maximize region coverage. Each attractor initiation state is validated to handle all goal object poses within its neighborhood. Online execution retrieves the precomputed path to the appropriate representative state and executes the behavior rollout, hence enabling fast online planning.
@@ V-B Baseline Methods
Online Planning Baseline: The fully online baseline represents the common approach for behavior-based manipulation planning. For each query with target object pose $w_{g}$, it first samples initiation states for the specified behavior $\sigma$ using the GetInitStates function, then attempts to generate collision-free motion plans to these states. Upon successful plan generation, it executes the behavior policy $\pi_{\sigma}$. This approach incurs expensive online computation as it performs sampling, motion planning, and behavior rollout for every query. Additionally, finding an initiation state does not mean that the policy execution will succeed. We implement this baseline using the BiTRRT [25] planner from OMPL.
@@ V-C Experimental Setup
We conducted 100 trials for the grasping domain and 60 trials for the insertion domain, with all plans starting from the robot home state $s_{home}$. The chosen goal states cover diverse regions of the task space to capture varying spatial localities and manipulation complexities, enabling comprehensive evaluation across different planning and behavior execution scenarios. A timeout of 5 seconds was set for all planners.
(a) Shelf Grasping Environment
(b) Charger Insertion Environment
(c) Success Rate
(d) Planning Time(ms)
Figure 3: Experimental comparison of B-CTMP against baseline methods on the two tasks. (3(a)) The shelf grasping environment, where a fixed-base robot must grasp a known target object at any location on the shelf. (3(b)) The charger insertion environment, where the robot must precisely insert a charger into a port located anywhere within the confined box, with pose uncertainty due to perception noise. (3(c)) Success rates for collision-free motion planning and subsequent behavior execution show that baseline methods exhibit 10-30% failure rates due to unsuccessful behavior rollouts, while B-CTMP maintains consistently high performance. (3(d)) Planning times demonstrate B-CTMP’s sub-millisecond online query performance through fast lookup operations, compared to expensive online computation required by baseline methods.
@@ V-D1 Performance Comparison
Figure 3 presents the success rate and planning time for B-CTMP compared to the baseline methods. In both experiments, our planner maintained a 100% end-to-end success rate, whereas the baselines failed either at finding valid collision-free paths to initiation states or during behavior execution from the path terminal states. Our approach also demonstrates sub-millisecond online query performance through fast lookup operations, compared to expensive online computation required by baseline methods. Additionally, B-CTMP achieves significant memory efficiency, with its offline storage requiring up to 21x less memory than Vanilla CTMP. This efficiency stems from B-CTMP’s decomposition strategy. Rather than storing neighborhoods in robot configuration or end-effector space like Vanilla CTMP, where physically proximate configurations may be partitioned into separate regions due to greedy unreachability, B-CTMP organizes neighborhoods in the object space using high coverage behavior initiation states. This object-centric representation enables more coherent neighborhood structures that better reflect the underlying task geometry.
V-D2 Failure Mode Analysis
@@
The results reveal three critical limitations of the baselines that B-CTMP addresses. First, discovering or specifying $\widetilde{\mathcal{S}}$ (a set of feasible initiation states with high coverage) is computationally expensive online and error-prone. This challenge is especially evident in the Vanilla CTMP baseline, which requires practitioners to manually tune preprocessed regions by estimating regions in task space $\mathcal{S}_{\text{init}}$ which contain good initiation states with high target pose coverage. This leads to frequent planning failures, since estimating this set is non-trivial and is behavior-dependent. In contrast, B-CTMP automatically discovers and selects a curated subset of initiation states during preprocessing, ensuring that stored states are both memory-efficient and support valid behavior rollouts. Second, the experiments expose the fundamental inadequacy of treating behavior validation as an afterthought. Baseline methods decouple motion planning from behavior execution, leading to failures from initiation states that are kinematically reachable but behaviorally invalid. For the insertion domain, regions such as shelf corners and side walls proved particularly challenging, as these locations often yield initiation states with poor manipulability or suboptimal approach angles for successful behavior execution. Since B-CTMP explicitly simulates behavior rollouts during preprocessing, it selects appropriate paths and initiations that promote success. Third, B-CTMP demonstrates flexibility in handling domain-specific constraints such as uncertainty requirements. While baselines cannot account for the requirements needed for corrective motions under perception noise, B-CTMP actively integrates manipulability constraints during preprocessing, ensuring selected initiation states maintain adequate dexterity throughout behavior execution.
@@
These results demonstrate that planning to kinematically feasible, collision-free initiation states alone is insufficient for reliable behavior-based manipulation. B-CTMP’s integrated approach of coupling motion planning with behavior validation during preprocessing achieves PR-Completeness by ensuring all stored states will result in successful behavior execution. This establishes both theoretical guarantees and practical advantages over existing methods that treat motion planning and behavior execution as separate, sequential processes.
VI Conclusion and Future Work

### v2

@@ Abstract
Recent progress in contact-rich robotic manipulation has been striking, yet most deployed systems remain confined to simple, scripted routines. One of the key barriers is the lack of motion planning algorithms that can provide verifiable guarantees for safety, efficiency and reliability. To address this, a family of algorithms called Constant-Time Motion Planning (CTMP) was introduced, which leverages a preprocessing phase to enable collision-free motion queries in a fixed, user-specified time budget (e.g., 10 milliseconds). However, existing CTMP methods do not explicitly incorporate the manipulation behaviors essential for object handling. To bridge this gap, we introduce the Behavioral Constant-Time Motion Planner (B-CTMP), an algorithm that extends CTMP to solve a broad class of two-step manipulation tasks: (1) a collision-free motion to a behavior initiation state, followed by (2) execution of a manipulation behavior (such as grasping or insertion) to reach the goal. By precomputing compact data structures, B-CTMP guarantees constant-time query in mere milliseconds while ensuring completeness and successful task execution over a specified set of states. We evaluate B-CTMP on two canonical manipulation tasks, shelf picking and plug insertion, in simulation and on a real robot. Our results show that B-CTMP unifies collision-free planning and object manipulation within a single constant-time framework, providing provable guarantees of speed and success for manipulation in semi-structured environments.
@@ I Introduction
Despite remarkable advances in robotic manipulation research, a persistent gap remains between the capabilities demonstrated in research laboratories and the requirements for real-world industrial deployment. This gap is especially pronounced in semi-static environments such as warehouse shelf picking (e.g., Amazon fulfillment centers [1]), bin picking in logistics, and precision assembly in manufacturing. A central challenge is the lack of verifiable guarantees on system performance—particularly safety, efficiency, and predictability—which are essential for practitioners in safety-critical and high-throughput applications.
@@ II-A Motion Planning with Preprocessing
To address these limitations, recent approaches focus on constructing collision-free regions in the configuration space rather than sampling discrete states, making motion within these regions computationally efficient during online planning. A prominent example of this region-based approach is Planning in Graphs of Convex Sets [15, 14] which decomposes the configuration space offline into collision-free convex sets [3, 18], enabling fast generation of smooth motion plans. These algorithms offer several advantages: they enable smooth trajectory generation through convex optimization and can handle complex geometric constraints naturally. However, they do not guarantee time bounds on online planning duration.
@@ III Preliminary
To build towards our aim, we first define coverage to capture behavior feasibility from specific robot states.
@@ IV-A Preprocessing Phase
Instead, we exploit a key insight: manipulation behaviors often exhibit spatial locality, where a single initiation state can cover multiple object configurations within a spatial region. For example, when grasping objects from a shelf, one pre-grasp pose (initiation state) may enable successful grasping of any object within a certain shelf region, as the behavior can adapt to the specific object pose during execution. Hence, leveraging this coverage property of initiation states, we can strategically select a reduced set of feasible initiation states $\widetilde{\mathcal{S}}\subseteq\displaystyle{\bigcup_{w_{g}\in\mathcal{G}}I_{\sigma}(w_{g})}$ such that: $\bigcup_{s_{i}\in\widetilde{\mathcal{S}}}n_{i}(s_{i})\supseteq\mathcal{G}$,which provides complete coverage over all desired goal states while requiring significantly fewer paths from $s_{home}$.
@@
Figure 2: Overview of the preprocessing phase. The planner computes a reduced set of feasible initiation states $\widetilde{\mathcal{S}}$, each connected to the robot’s home state $s_{home}$ by a precomputed path. Each $s_{i}\in\widetilde{\mathcal{S}}$ defines a neighborhood $n_{i}(s_{i})$ within the region of interest $\mathcal{G}$ of object states $w$ that can be reached through behavior execution. This compression is captured by attractor tuples, each representing an initiation state, its reachable neighborhood, and the stored path that enables efficient plan retrieval during online execution.
@@
Once the algorithm encounters an object state that cannot be reached by the behavior rollout from $s_{i}$, that state is recorded and subsequently used to seed coverage for a different local region of interest. After all the candidate neighborhoods are expanded, the largest neighborhood (measured in terms of the maximum distance $r$ between the attractor object state $w_{attr}$ and the closest frontier state) is selected to get added to the stored library, and the initiation state corresponding to the neighborhood is chosen as the
@@ V Experiments
We evaluate our B-CTMP on canonical manipulation tasks across two environments–shelf picking and plug insertion, in both simulation (UR10e robot and Yaskawa hc10dtp) and physical (UR10e robot) setups.
@@ V-A1 Shelf Grasping
For our first task, we consider a shelf-picking scenario commonly encountered in industrial warehouse automation, where a robotic manipulator must retrieve objects from structured storage environments. The manipulation task involves grasping a target object whose pose is estimated from perception and specified in the world coordinate frame by $w_{g}\in\mathrm{SE}(3)$.
Grasp Behavior: The grasp behavior $\sigma_{grasp}$ depends on two key points: the pre-grasp pose (an initiation state) from which we execute the policy, and the grasp pose, where we activate a closure sequence to grasp the object. Since the object model is known, both poses are computed from the estimated object pose $w_{g}$, meaning the behavior must handle uncertainty introduced by perception noise in the pose estimate. The behavior policy $\pi_{\sigma_{grasp}}$ uses Jacobian control to gradually move the end-effector through differential kinematics between these points, enabling smooth grasp execution. To ensure robustness under perception uncertainty, we impose a manipulability constraint during policy execution. We define the minimum manipulability radius as $r_{\min}(q)=\min_{i}\sqrt{\lambda_{i}\!\big(J(q)J(q)^{\mathsf{T}}\big)}$, where the manipulability ellipsoid radii are derived from the Jacobian eigenvalues.
Behavior Implementation: The GetInitStates function for the grasp behavior outputs a subset of valid pre-grasp poses (initiation states), computed as follows. First, we compute candidate grasp poses with valid IK solutions using the known object model. Local surface normals are then used to calculate antipodal scores that quantify grasp stability by measuring alignment between gripper fingers and object surface. Using the antipodal scores as grasp ranking heuristics, we select the top $K$ grasps to attempt behavior execution from. The corresponding pre-grasp poses are computed by applying a fixed retraction transformation along each grasp’s approach direction.
@@ V-A2 Charger Insertion
The second task we consider is precision charger insertion, commonly required in automated charging systems and electronic assembly. This task requires inserting a charger connector into a target port located at $w_{g}\in SE(3)$. To complete the task successfully, the planner must navigate strict geometric constraints along with uncertainty in the perceived port pose.
Insertion Behavior: Since the port geometry is known and uniquely dictates a single insertion pose, the behavior only requires the initiation state from which to rollout the policy. As in the grasping task, due to the high precision required and perception uncertainty in $w_{g}$, we rely on visual-based Jacobian control for our policy $\pi_{\sigma_{insert}}$ and impose the same manipulability constraint to ensure robust execution.
Behavior Implementation: We compute the required insertion pose from the known port geometry and estimated pose $w_{g}$, then apply a predefined transformation to obtain the corresponding pre-insert initiation state, which is output by the GetInitStates function and subsequently used for behavior rollout and region expansion.
V-A3 B-CTMP Integration
In both tasks, B-CTMP leverages the behavior during offline preprocessing to compute a compact set of representative initiation states with their corresponding neighborhoods. It does so by imposing the manipulability constraint during behavior rollouts, marking a rollout as infeasible if $r_{\min}(q)<\epsilon$, where $\epsilon$ represents the perception noise bound [^1] The perception noise bound $\epsilon$ is determined from the camera calibration process, which provides an upper bound on the pose estimation error.. This ensures that selected attractor pre-grasp or pre-insert poses maintain adequate dexterity for corrective motions across their neighborhoods, making them capable of handling many different target object configurations—including those subject to perception-induced pose uncertainty. During online execution, the precomputed path to the appropriate representative initiation state is retrieved and the behavior rollout is executed, enabling fast online planning.
@@ V-B Baseline Methods
Online Planning Baseline: The fully online baseline represents the common approach for behavior-based manipulation planning. For each query with target object pose $w_{g}$, it first samples initiation states for the specified behavior $\sigma$ using the GetInitStates function, then attempts to generate collision-free motion plans to these states. Once a valid plan is found, it executes the behavior policy $\pi_{\sigma}$ from the terminal state of that plan. We additionally evaluate a ”re-planning” variant, where instead of committing to the first feasible motion plan, the planner iterates over multiple sampled initiation states, repeating the attempted motion planning and policy rollout process.
This approach incurs expensive online computation as it performs sampling, motion planning, behavior rollout, and optionally replanning for every query. Additionally, finding an initiation state does not mean that the policy execution will succeed. We implement this baseline using the BiTRRT [4] planner from OMPL.
@@ V-C Experimental Setup
Our experimental evaluations consist of 100 grasping and 60 insertion trials in simulation, alongside 50 physical robot trials for each task. In all experiments, plans start from the robot home state $s_{home}$. The chosen goal states cover diverse regions of the task space to capture varying spatial localities and manipulation complexities, enabling comprehensive evaluation across different planning and behavior execution scenarios. A timeout of 5 seconds was set for all planners.
Grasping Object from a Shelf
[…]
Feasible Task
Execution Success [%]
Infeasible Task [%]
Online Planning
Time [ms]
100 18 1.8
[…]
Goal Region
Volume $cm^{3}$]
Preprocessing
Time [hours]
Memory
Reduction [%]
1260 0.28 99.6
2500 5.90 94.0
5600 14.5 92.7
Plug Insertion
[…]
Feasible Task
Execution Success [%]
Infeasible Task [%]
Online Planning
Time [ms]
100 20 1.3
[…]
Goal Region
Volume $cm^{3}$]
Preprocessing
Time [mins]
Memory
Reduction [%]
21000 5.46 56.0
45000 9.94 61.8
96000 26.9 67.1
Figure 3: Real-robot experiments evaluating B-CTMP across two manipulation tasks (50 trials per task): shelf grasping task (top), and charger insertion task (bottom). Each table reports: (i) the online execution success rate over feasible task instances, (ii) the percentage of infeasible instances correctly identified as infeasible (computed relative to the total number of trials), and (iii) the online planning time. We additionally present representative infeasible cases. For grasping (top), the object is placed at a location that becomes unreachable due to collisions with static obstacles across all possible grasp poses. For insertion (bottom), certain object poses induce joint singularities, resulting in loss of manipulability and failure to reach the target port. Red bounding boxes indicate the specific affected joints in each case. Finally, we report preprocessing time and the memory reduction achieved relative to a naive baseline that stores the full, uncompressed precomputed data.
(a) Shelf Grasping Task- Success Rate and Planning Time
(b) Charger Insertion Task- Success Rate and Planning Time
Figure 4: Success rates for collision-free motion planning and subsequent behavior execution show that baseline methods exhibit 10-30% failure rates due to unsuccessful behavior rollouts and initiations, while B-CTMP maintains consistently high performance. Planning times demonstrate B-CTMP’s sub-millisecond online query performance, compared to expensive online computation required by baselines.
@@ V-D1 Simulation Experiments
Figure 4 presents the success rate and planning time for B-CTMP compared to the baseline methods in simulation. In both experiments, our planner maintained a 100% end-to-end success rate, whereas the baselines failed either at finding valid collision-free paths to initiation states or during behavior execution from the path terminal states. Our approach also demonstrates sub-millisecond online query performance through fast lookup operations, compared to expensive online computation required by baseline methods. This efficiency stems from B-CTMP’s decomposition strategy. Rather than storing neighborhoods in robot configuration or end-effector space like Vanilla CTMP, where physically proximate configurations may be partitioned into separate regions due to greedy unreachability, B-CTMP organizes neighborhoods in the object space using high coverage behavior initiation states. This object-centric representation enables more coherent neighborhood structures that better reflect the underlying task geometry.
V-D2 Physical Experiments
Figure 3 presents the real-world experimental setup and results for the grasping and insertion tasks. In both domains, B-CTMP maintains a 100% execution success rate despite perception noise and object pose randomization. We additionally report the percentage of task-infeasible instances — cases where a sampled object location lies within the goal region yet is invalid and the task cannot be accomplished. Such infeasibility arises when the task requires motions that are kinematically infeasible due to collisions or joint singularities. To validate this, we attempted execution at several such locations and consistently observed singularity failures or collisions, confirming both the infeasibility of these configurations and the correctness of the preprocessed feasibility regions. Notably, B-CTMP reports task infeasibility in constant time, without requiring any additional planning or execution attempts. The accompanying video further demonstrates our approach, including demonstrations of infeasible planning instances.
V-D3 Preprocessing Analysis
The tables in Figure 3 report preprocessing time and memory reduction across goal region volumes for each task. Memory reduction is measured relative to a naive baseline that explicitly computes and stores individual paths to a feasible behavior initiation state for every discretized object pose within the goal region. B-CTMP achieves over 90% memory reduction in the grasping task and over 55% in the insertion task.This reduction has direct implications for online planning efficiency, as plan retrieval time scales with cache size and excessive memory overhead can introduce latency during real-time execution.
We note that preprocessing time and memory requirements depend on factors beyond goal region volume. For example, we see that preprocessing for the grasping task takes significantly longer than for the insertion task, because the latter has only one unique initiation state per object pose and hence fewer required rollout and planning attempts. Moreover, regions near the robot’s kinematic limits or in cluttered areas may require finer discretizations, and some behaviors may have sparse initiation states with low coverage. Consequently, these statistics vary with goal region geometry, skill definition and sensitivity, and object-pose resolution.
V-D4 Baseline Failure Mode Analysis
@@
The results reveal three critical limitations of the baselines that B-CTMP addresses. First, identifying $\widetilde{\mathcal{S}}$ (a set of feasible initiation states with high coverage) is computationally expensive online and error-prone, especially in the Vanilla CTMP baseline, which requires practitioners to manually tune preprocessed regions by estimating regions in task space $\mathcal{S}_{\text{init}}$. In contrast, B-CTMP automatically curates a subset of initiation states during preprocessing, ensuring that stored states are both memory-efficient and support valid behavior rollouts. Second, the experiments expose the fundamental inadequacy of treating behavior validation as an afterthought. Baseline methods decouple motion planning from behavior execution, leading to failures from initiation states that are kinematically reachable but behaviorally invalid. For the insertion domain, regions such as shelf corners and side walls proved particularly challenging, as these locations often yield initiation states with poor manipulability or suboptimal approach angles for successful behavior execution. Since B-CTMP explicitly simulates behavior rollouts during preprocessing, it selects appropriate paths and initiations that promote success. Third, B-CTMP flexibly incorporates domain-specific constraints such as uncertainty requirements. While baselines do not account for the requirements needed for corrective motions under perception noise, B-CTMP actively reasons about manipulability constraints, ensuring selected initiation states maintain adequate dexterity throughout behavior execution.
@@
These results demonstrate the benefit of B-CTMP’s integrated approach of coupling motion planning with behavior validation during preprocessing. The algorithm achieves PR-Completeness by ensuring all stored states will result in successful behavior execution. This establishes both theoretical guarantees and practical advantages over existing methods that treat motion planning and behavior execution as separate, sequential processes.
VI Limitations
We acknowledge several limitations of our proposed B-CTMP approach. First, our algorithm relies on prior knowledge of the workspace geometry and the goal region. While this can be restrictive in dynamic environments, we note that in the semi-static settings targeted by this paper, the environment layout is typically known a priori and remains largely fixed. Hence, the assumption aligns with real deployment conditions. Second, B-CTMP assumes access to a high-fidelity behavior simulator to identify high-coverage initiation states and support execution guarantees and PR-completeness. For closed-loop Jacobian-based control, as considered in this work, this is reasonable due to deterministic feedback-driven state evolution. However, for some learned behaviors, model uncertainty and sim2real gap can be a barrier to offline validation. Extending B-CTMP to such behaviors remains important future work.
VII Conclusion